Prove or disprove the following:

If f(n) =O(g(n)) then nf(n) = O(ng(n))
Answer #1

Since f(n) = O(f(n)), this means that as per Big-O definition, there exists a positive constant c such that for n >= n0, we have

0 <= f(n) <= c*g(n)

Raising to the power of n, we get

0 <= n^f(n) <= n^(c * g(n))

=> 0 <= n^f(n) <= n^(c * g(n))

And, as per Big-O definition,

nf(n) = O(ng(n))